OpenStudy (anonymous):

Let f(x) = (11 ln x)/(8 x). what is the derivative? should look like this: f' = (11 + A ln x)/(B x^C)

OpenStudy (amistre64):

its whatever the quotient rule pops out id asume

OpenStudy (amistre64):

you could put the 11/8 to the side and work the rest

OpenStudy (amistre64):

\[\frac{a}{b}=ab^{-1}\]so you could rewrite it as a product rule

OpenStudy (anonymous):

\[f(x)=\frac{11\ln x}{8x}=(11\ln x)(8x)^{-1}.\]If you are lazy, like me, you can apply the product rule instead.

OpenStudy (amistre64):

mathikers are quite lazy ;)

OpenStudy (anonymous):

wait so (11lnx)(8x)^-1 then simplify?

OpenStudy (amistre64):

take the derivative of that using the product rule and then simplify; yes